### 20 Apr 2026 v2

Abstract
Non-commutative electrodynamics obtained through the Seiberg–Witten map ceases to have equivalent action-level and equation-level realizations once fixed external currents are introduced, and in the action-level construction associated with the Banerjee current map the canonical location of this source-induced obstruction has remained unclear. Working in the full phase space and treating the current as prescribed and non-dynamical, we apply the Dirac–Bergmann algorithm without imposing current conservation as an external condition. The preservation of the Gauss-type secondary constraint produces a third-stage candidate whose phase-space expression is shown to be algebraically identical, at first order in the non-commutativity parameter and for purely space–space non-commutativity, to the canonical pullback of the divergence of the mapped Euler–Lagrange equations. This identity locates the source-compatibility obstruction directly within the Dirac chain. For generic inhomogeneous sources, the next consistency step feeds this object back into the primary multiplier through a source-dependent kernel, so the chain closes by multiplier fixing rather than by the generic appearance of a quaternary constraint. Reduced-phase-space results, including the gauge generator, Dirac brackets and degree-of-freedom count, are obtained only in a restricted sufficient first-class subcase; no broader claim is made for arbitrary source profiles.
keywords:
Non-commutative electrodynamics, Seiberg–Witten map, Dirac–Bergmann analysis, External currents, Constraint structure, Gauge obstruction
1 Introduction
Non-commutative (NC) field theory deforms the coordinate algebra according to
$\bigl[x^{\mu},x^{\nu}\bigr]=i\,\theta^{\mu\nu},$ (1.1)
where $\theta^{\mu\nu}$ is a constant antisymmetric tensor [1, 2]. Through the Seiberg–Witten (SW) map, NC gauge theories can be rewritten order by order in $\theta^{\mu\nu}$ in terms of ordinary fields and standard spacetime coordinates [3, 4, 5, 6, 7, 8, 9]. In the absence of sources, the first-order map leads to the familiar gauge-invariant weakly nonlinear deformation of Maxwell theory [10, 11].
With external sources the situation is subtler. In ordinary Maxwell theory a prescribed current is compatible with gauge symmetry only if it is conserved. In the NC setting, Adorno et al. [12] showed that the action-level and equation-level implementations of the SW map are no longer equivalent in the presence of fixed external currents: mapping after deriving the equations preserves gauge covariance, whereas mapping the action before variation does not. The Hamiltonian meaning of that mismatch, and in particular the point at which it enters the Dirac chain, has remained less clear.
Related external-source analyses show that the nature of the source itself is part of the problem rather than a secondary modeling choice. Cabo and Shabad treat a non-Abelian classical current under the restriction $J_{4}^{\,a}=0$, so that the source modifies the canonical consistency conditions without producing the same temporal shift of Gauss’s law examined here [13]. Przeszowski’s canonical path-integral analysis of Yang–Mills theory with arbitrary external sources goes further in a different direction: the naive source coupling is replaced by a modified source sector before quantization, precisely because the canonical structure is sensitive to the way the external current is introduced [14]. Sikivie and Weiss, by contrast, treat static external sources as fixed backgrounds used to define distinct classical solution sectors, such as Coulomb and screening branches, rather than to analyze first-/second-class structure in the sense of Dirac–Bergmann [15]. In NC Maxwell theory, Kruglov adds a prescribed Abelian current directly to the first-order SW-expanded Lagrangian and studies the resulting nonlinear field equations and energy–momentum tensors, but without a Hamiltonian source-sector analysis of the type carried out here [16]. These comparisons delimit the present problem more sharply: the issue is not merely that an external source is present, but how a fixed four-current enters the canonical consistency chain of the action-level SW-mapped theory.
To date, no Dirac–Bergmann analysis of the first-order action-level SW-mapped theory with fixed external currents has been carried out: the existing treatments either operate exclusively at the Lagrangian level [12, 16], or address external sources in canonical frameworks that do not involve the SW map [13, 14]. As a consequence, the precise point in the canonical consistency chain at which the source-induced obstruction enters—and the way in which it depends on the choice of SW current map—has remained undetermined.
This paper fills that gap for the first-order action-level realization based on the Banerjee current map [17]. We treat $J^{\mu}$ as a prescribed non-dynamical source and apply the Dirac–Bergmann algorithm in the full phase space, keeping $A_{0}$ and $\pi^{0}$ as canonical variables throughout. The analysis is restricted to first order in $\theta$ and to purely spatial non-commutativity. Within that framework, the secondary Gauss-type constraint already shows where the source enters the canonical structure, while its preservation generates a third-stage source-dependent object whose canonical form is shown independently to be algebraically identical to the canonical pullback of the divergence of the Banerjee Euler–Lagrange equations. Because this identity is tied to the specific $F_{\alpha\beta}$-dependent terms of the Banerjee map, it is not a generic consequence of any action-level SW realization; rather, it characterizes the map dependence of the obstruction at the canonical level. The next consistency step closes the generic source sector by multiplier fixing—itself a positive structural result, not merely a limitation—whereas an exact reduced Hamiltonian description exists only in a restricted sufficient first-class subcase.
The rest of the paper is organized as follows. Section 2 fixes the working model: the conventions (Sec. 2.1), the source-free SW-mapped action in closed commutative form (Sec. 2.2), and the current-sector choice—the Banerjee map—that differentiates the present construction from that of Adorno et al. (Sec. 2.3). Section 3 then develops the Dirac–Bergmann analysis of the unrestricted external-current problem, up to and including the tertiary stage, and establishes the central identity of the paper (Proposition 1): the tertiary-stage candidate coincides, in the canonical framework, with the divergence of the mapped Euler–Lagrange equations. Section 4 develops the reduced phase-space description that becomes available only in a restricted sufficient first-class subcase, and uses it to compare the Banerjee reduction with the gauge-invariant equation-level formulation of Ref. [12]. The structural consequences, scope, and limitations are analyzed in Sec. 5, and the findings are summarized in Sec. 6.
2 Working model and current map
2.1 Conventions and notation
We first fix the notation needed in the rest of the paper. We work in Minkowski spacetime with metric signature $\eta^{\mu\nu}=\mathrm{diag}(+1,-1,-1,-1)$ and set $\hbar=c=1$. Greek indices $\mu,\nu,\rho,\ldots$ run over $\{0,1,2,3\}$, while Latin indices $i,j,k,\ldots$ are restricted to spatial components $\{1,2,3\}$. The non-commutativity parameter $\theta^{\mu\nu}=-\theta^{\nu\mu}$ is a constant antisymmetric tensor with dimensions $\text{length}]^{2}$.
To avoid issues with unitarity and acausal propagation, we consider purely space-space non-commutativity and set $\theta^{0i}=0$ identically. From the spatial components $\theta^{ij}$ we form the NC vector
$\theta^{i}\equiv\frac{1}{2}\,\varepsilon^{ijk}\,\theta_{jk}\,$, (2.1)
where $\varepsilon^{ijk}$ is the Levi-Civita symbol, normalized by $\varepsilon^{123}=+1$. The commutative electromagnetic field strength is $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, and we decompose it into electric and magnetic components as
$E^{i}=-F^{0i}\,,\quad B^{i}=\frac{1}{2}\,\varepsilon^{ijk}F_{jk}\,.$ (2.2)
2.2 Source-free mapped action
We begin with the source-free NC Maxwell action, since the source sector will later be added on top of the same mapped gauge block [17, 18]
$S_{\text{NC}}=\int d^{4}x\,\left(-\frac{1}{4}\,\hat{F}_{\mu\nu}\star\hat{F}^{\mu\nu}\right),$ (2.3)
where the NC field strength is given by [^1] In order to construct the non-commutative version of a given field theory, we replace the ordinary pointwise product of fields in the action by the Moyal $\star$-product. The only identities needed below are the standard first-order expansion, cyclicity of the integral under suitable boundary conditions, and the Leibniz rule for ordinary derivatives.
$\hat{F}_{\mu\nu}={}$ $\,\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$ (2.4)
$\,-i[\hat{A}_{\mu},\hat{A}_{\nu}]_{\star}\,.$
For NC $U(1)$, the ordinary commutator vanishes, but the $\star$-commutator yields derivative corrections starting at $O(\theta)$; these are consistently captured in the SW expansion. The Seiberg–Witten map expresses the NC quantities $(\hat{A}_{\mu},\hat{F}_{\mu\nu})$ in terms of their commutative counterparts $(A_{\mu},F_{\mu\nu})$ as a power series in $\theta$ [5, 6]. To first order one finds
$\hat{A}_{\mu}={}$ $\,A_{\mu}-\frac{1}{2}\,\theta^{\alpha\beta}\,A_{\alpha}\left(\partial_{\beta}A_{\mu}+F_{\beta\mu}\right)$ (2.5)
$\,+\mathcal{O}(\theta^{2})\,.$
From this, the induced map for the field strength follows:
$\hat{F}_{\mu\nu}={}$ $\,F_{\mu\nu}+\theta^{\alpha\beta}\Bigl(F_{\mu\alpha}F_{\nu\beta}-A_{\alpha}\partial_{\beta}F_{\mu\nu}\Bigr)$ (2.6)
$\,+\mathcal{O}(\theta^{2})\,.$
Substituting Eq. (2.6) into Eq. (2.3) and using the standard cyclicity of the $\star$-product under the integral, we may discard total derivatives. The resulting Lagrangian can be written entirely in terms of the commutative field strength $F_{\mu\nu}$:
$\mathcal{L}_{\text{ws}}=$ $-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\frac{1}{8}\,\theta^{\alpha\beta}\,F_{\alpha\beta}\,F_{\mu\nu}F^{\mu\nu}$
$-\frac{1}{2}\,\theta^{\alpha\beta}\,F_{\mu\alpha}\,F_{\nu\beta}\,F^{\mu\nu}\,.$ (2.7)
Relative to the standard Maxwell Lagrangian, NC electrodynamics acquires terms cubic in $F_{\mu\nu}$, signalling that non-commutativity induces non-linearities already at the classical level.
Using Eqs. (2.2) and (2.1), the Lagrangian can be rewritten in the more transparent form
$\mathcal{L}_{\text{ws}}$ $=\tfrac{1}{2}(E^{2}-B^{2})\bigl[1+(\boldsymbol{\theta}\!\cdot\!\mathbf{B})\bigr]$
$\quad-(\boldsymbol{\theta}\!\cdot\!\mathbf{E})(\mathbf{E}\!\cdot\!\mathbf{B})+\mathcal{O}(\theta^{2})\,.$ (2.8)
This expression makes explicit how NC corrections couple the vector $\boldsymbol{\theta}$ to the magnetic field and generate a mixed $\mathbf{E}$$\mathbf{B}$ term that vanishes in the commutative limit $\theta\to 0$. As we will see in Sec. 3, these terms have direct implications for the constraint structure, specifically for the form of the secondary (Gauss-type) constraint and the gauge-breaking mechanism that emerges when external sources are included.
2.3 The Banerjee current map and relation to earlier work
Before turning to the constraint analysis, we specify the current-sector choice and its relation to earlier work. In the absence of external sources, both the Banerjee and the Adorno et al. constructions reduce to the same first-order SW-mapped Maxwell theory: with the convention matching $16\pi c=4$, $g=1$, the source-free mapped action of Ref. [12] reduces exactly to $\mathcal{L}_{\text{ws}}$ in Eq. (2.7), with $f_{\mu\nu}\equiv F_{\mu\nu}$. The source-free sector is therefore common to both formulations; the difference arises when a fixed external current is included.
For the current sector we adopt the standard first-order Banerjee map [17], corresponding to the parameter choice $(c_{1},c_{2},c_{3})=(1,\tfrac{1}{2},0)$:
$\hat{J}^{\mu}$ $=J^{\mu}-A_{\alpha}\,\theta^{\alpha\beta}\,\partial_{\beta}J^{\mu}+F_{\alpha\beta}\,J^{\beta}\,\theta^{\mu\alpha}$
$\quad+\frac{1}{2}F_{\alpha\beta}J^{\mu}\theta^{\alpha\beta}+\mathcal{O}(\theta^{2})\,.$ (2.9)
This parameter choice satisfies the SW consistency condition $\delta_{\lambda}\hat{J}^{\mu}=0$ for ordinary gauge transformations $\delta_{\lambda}A_{\mu}=\partial_{\mu}\lambda$, as demonstrated in Ref. [17] (see also [21] for the general analysis of SW-map ambiguities in the current sector and [23] for the implications of these maps in the anomalous current algebra). The ambiguity at first order in $\theta$ resides in the relative weight of the $F_{\alpha\beta}$-dependent terms, which is fixed by the condition that the mapped current transforms covariantly; the choice $(c_{1},c_{2},c_{3})=(1,\tfrac{1}{2},0)$ is the unique solution within the parametrization of Ref. [17].
Relative to the minimal current map used by Adorno et al.,
$\check{j}^{\mu}=j^{\mu}+g\,\theta^{\alpha\beta}A_{\alpha}\partial_{\beta}j^{\mu}\,$, (2.10)
the Banerjee choice (2.9) contains additional $F_{\alpha\beta}$-dependent terms. These terms belong to the first-order SW-map ambiguity of the current sector and are consistent with the analysis of Ref. [12]. They leave the source-free theory unchanged but modify the source coupling at $\mathcal{O}(\theta)$, which is precisely the effect relevant for the Hamiltonian constraint structure analyzed below.
Adorno et al. showed that applying the SW map at the action level before variation yields equations of motion that are not gauge covariant when $J^{\mu}$ is prescribed and non-dynamical, whereas applying the map after variation (at the equation-of-motion level) preserves gauge covariance. They conclude that the consistent route in the presence of external sources is the equation-level one. The question addressed in the present work is therefore not whether the SW-mapped action is gauge covariant—that it is not, as Adorno et al. established—but whether its constrained Hamiltonian structure provides a precise canonical diagnosis of that obstruction, and how the diagnosis depends on the choice of SW current map. From the viewpoint of constrained dynamics, the Lagrangian mismatch of Ref. [12] becomes a question about closure, preservation, and first-/second-class classification in the full Dirac–Bergmann [19, 20] framework, which is developed in Sec. 3.
3 Dirac–Bergmann analysis in full phase space
With the working Lagrangian $\mathcal{L}=\mathcal{L}_{\rm ws}+\mathcal{L}_{J}$ and the Banerjee current map fixed in Sec. 2, we now submit the theory to the Dirac–Bergmann algorithm in the full phase space, treating $J^{\mu}$ as prescribed and non-dynamical and without imposing $\partial_{\mu}J^{\mu}=0$ from the outside. The goal of this section is twofold. First, to identify the canonical locus at which the external source first enters the consistency chain, which will turn out to be the tertiary stage (Sec. 3.3, Proposition 1). Second, to characterize the algebraic structure of that tertiary obstruction through the three source kernels $\Theta^{l}$, $\Sigma^{lk}$, $\Xi^{l}$, which determine how the algorithm closes and in which restricted sector a standard first-class reduction remains available (Sec. 3.5). Throughout this section, no restriction is placed on the source profile; the sectorial discussion emerges from the algorithm itself.
3.1 Source Lagrangian, primary constraint, and primary Hamiltonian
We split the working Lagrangian into a source-free gauge block and a mapped source block,
$\mathcal{L}=\mathcal{L}_{\text{ws}}+\mathcal{L}_{J}\,$, (3.1)
where $\mathcal{L}_{\text{ws}}$ is the source-free piece in Eq. (2.8), and $\mathcal{L}_{J}$ denotes the coupling to the prescribed external current after the SW map. Substituting the Banerjee current map (2.9) into the interaction term and using the standard properties of the $\star$-product, one obtains
$\begin{split}\mathcal{L}_{J}&=-A_{\mu}\Bigl[J^{\mu}+\theta^{\mu\alpha}J^{\beta}\Bigl(\partial_{\alpha}A_{\beta}-\tfrac{1}{2}\partial_{\beta}A_{\alpha}\Bigr)\Bigr]\\ &\qquad+\mathcal{O}(\theta^{2})\,.\end{split}$ (3.2)
With the source Lagrangian in hand, we now proceed to the canonical formulation. The momenta conjugate to $A_{\lambda}$ are defined as
$\pi^{\lambda}\equiv\frac{\delta\mathcal{L}}{\delta\dot{A}_{\lambda}}\,$, (3.3)
where the dot denotes $\partial_{0}$. Since $\mathcal{L}$ does not contain $\dot{A}_{0}$, we immediately obtain the primary constraint
$\phi_{1}\equiv\pi^{0}\approx 0\,.$ (3.4)
The spatial components of the momentum decompose into source-free and source-dependent contributions,
$\pi^{i}=[\pi^{i}]_{\text{ws}}+[\pi^{i}]_{J}\,$, (3.5)
where
$[\pi^{i}]_{\text{ws}}$ $=\bigl[\tfrac{1}{2}F_{kl}\theta^{kl}-1\bigr]F^{0i}+F^{0j}F_{jk}\theta^{ki}$
$\quad+F^{0m}F_{kj}\eta_{lm}\eta^{ji}\theta^{lk}\,$, (3.6)
$[\pi^{i}]_{J}$ $=\tfrac{1}{2}\theta^{ji}A_{j}J^{0}\,.$ (3.7)
The combined spatial momentum is therefore
$\pi^{i}$ $=\bigl[\tfrac{1}{2}F_{kl}\theta^{kl}-1\bigr]F^{0i}+F^{0j}F_{jk}\theta^{ki}$
$\quad+F^{0m}F_{kj}\eta_{lm}\eta^{ji}\theta^{lk}+\tfrac{1}{2}\theta^{ji}A_{j}J^{0}\,.$ (3.8)
It will be useful throughout the paper to isolate the purely field-strength, source-independent $\mathcal{O}(\theta)$ deformation of the momentum. Writing
$\pi^{i}=-\,F^{0i}+\mathcal{K}^{i}(F)+\tfrac{1}{2}\theta^{ji}A_{j}J^{0},$ (3.9)
the field-strength bilinear kernel is
$\mathcal{K}^{i}(F)={}$ $\tfrac{1}{2}\,(\theta^{kl}F_{kl})\,F^{0i}+\theta^{ki}F^{0j}F_{jk}$ (3.10)
$+\theta^{lk}\eta_{lm}\eta^{ji}F^{0m}F_{kj}\,.$
Thus $\mathcal{K}^{i}(F)$ collects all the $\mathcal{O}(\theta)$ field-strength bilinear corrections that deform the Gauss-type canonical structure without introducing explicit dependence on the potential beyond the separate source term $\tfrac{1}{2}\theta^{ji}A_{j}J^{0}$. This decomposition will be used both in the compact form of the tertiary-stage analysis (Sec. 3.3) and in the on-shell Gauss-law comparison (Sec. 4.5).
Solving Eq. (3.8) perturbatively for $F^{0i}$ to first order in $\theta$ via $M^{i}{}_{j}=\delta^{i}{}_{j}+\mathcal{O}(\theta)$, one finds
$F^{0i}$ $=\pi^{i}\Bigl(\tfrac{1}{2}F_{kl}\theta^{lk}-1\Bigr)-\pi^{k}\bigl(F_{kl}\theta^{li}+F_{ml}\eta^{im}\eta_{nk}\theta^{ln}\bigr)$
$\quad+[\pi^{i}]_{J}\,.$ (3.11)
The canonical Hamiltonian splits into a source-free and a source-dependent part,
$H_{c}=H_{\text{ws}}+H_{J}=\int\left(\mathcal{H}_{\text{ws}}+\mathcal{H}_{J}\right)d^{3}x\,$, (3.12)
with
$\mathcal{H}_{\text{ws}}$ $=\pi_{i}\partial^{i}A_{0}+\tfrac{1}{4}F_{ij}F^{ij}\Bigl[1-\tfrac{1}{2}F_{ml}\theta^{ml}\Bigr]$
$\quad-\tfrac{1}{2}\eta_{00}\eta_{ij}\pi^{i}\pi^{j}\Bigl[1+\tfrac{1}{2}F_{ml}\theta^{ml}\Bigr]$
$\quad+\tfrac{1}{2}F^{ip}F_{li}F_{pm}\theta^{ml}$
$\quad-F_{li}\eta_{00}\eta_{mj}\pi^{i}\pi^{j}\theta^{ml}\,$, (3.13)
and
$\mathcal{H}_{J}$ $=A_{0}J^{0}+A_{j}J^{j}$
$\quad+\tfrac{1}{2}\theta^{jl}\bigl[A_{j}J^{0}\partial_{l}A_{0}+A_{j}J^{k}\partial_{l}A_{k}$
$\quad\qquad+A_{j}J^{0}\eta_{00}\eta_{li}\pi^{i}-A_{j}F_{kl}J^{k}\bigr]\,.$ (3.14)
The primary Hamiltonian is then
$H_{1}=\int\left(\mathcal{H}_{\text{ws}}+\mathcal{H}_{J}+u_{1}\,\phi_{1}\right)d^{3}x\,$, (3.15)
where $u_{1}$ is a Lagrange multiplier enforcing the primary constraint.
3.2 Secondary constraint and the tertiary-stage candidate
With the primary constrained surface defined and the primary Hamiltonian constructed, the Dirac–Bergmann algorithm requires us to demand the time preservation of the primary constraint, $\dot{\phi}_{1}\approx 0$. To evaluate this and the subsequent consistency conditions, we use the fundamental equal-time Poisson brackets
$\{A_{\mu}(x),\pi^{\nu}(y)\}=\delta^{\nu}_{\mu}\,\delta^{3}(x-y),$ (3.16)
and
$\{F_{ij}(x),\pi_{k}(y)\}=\delta^{k}_{i}\,\partial_{j}\delta^{3}(x-y)-\delta^{k}_{j}\,\partial_{i}\delta^{3}(x-y).$ (3.17)
Using Eq. (3.4) and the explicit form of Eq. (3.15), the condition $\{\phi_{1},H_{1}\}=0$ yields the secondary constraint
$\phi_{2}=\partial_{i}\pi^{i}-J^{0}+\frac{1}{2}\,\theta^{kl}\,\partial_{l}(A_{k}J^{0})\approx 0.$ (3.18)
In the commutative limit, this reduces to Gauss’s law, $\partial_{i}\pi^{i}=J^{0}$.
To test whether the algorithm closes or generates further constraints, we must also preserve $\phi_{2}$ in time. Because the external current $J^{\mu}(x)$ is a prescribed function (not a dynamical variable), its time derivative contributes through an explicit partial derivative. Hence
$\dot{\phi}_{2}=\{\phi_{2},H_{c}\}_{\text{PB}}+\frac{\partial\phi_{2}}{\partial t}\bigg|_{\rm expl},$ (3.19)
with
$\frac{\partial\phi_{2}}{\partial t}\bigg|_{\rm expl}=-\partial_{0}J^{0}+\frac{1}{2}\theta^{kl}\partial_{l}\!\left(A_{k}\,\partial_{0}J^{0}\right).$ (3.20)
The canonical Hamiltonian splits as $H_{c}=H_{\text{ws}}+H_{J}$ (source-free and source parts), so the bracket decomposes accordingly:
$\{\phi_{2},H_{c}\}_{\text{PB}}=\{\phi_{2},H_{\text{ws}}\}_{\text{PB}}+\{\phi_{2},H_{J}\}_{\text{PB}}.$ (3.21)
Using the functional-derivative form of the Poisson bracket, with the intermediate contractions collected in Appendix A, one finds
$\{\phi_{2},H_{\text{ws}}\}{}$ $=\frac{1}{2}\theta^{kl}(\partial_{l}J^{0})\,\partial_{k}A_{0}$ (3.22)
$\quad-\frac{1}{2}\theta^{jl}\eta_{00}\eta_{jn}\,\partial_{l}\!\bigl(J^{0}\pi^{n}\bigr),$
and
$\{\phi_{2},H_{J}\}{}$ $=-\partial_{i}J^{i}-\frac{1}{2}\theta^{il}\partial_{i}(J^{0}\partial_{l}A_{0})$
$\quad+\frac{1}{2}\theta^{jl}\partial_{i}\partial_{l}(J^{i}A_{j})-\frac{1}{2}\theta^{il}\partial_{i}(J^{k}\partial_{l}A_{k})$
$\quad-\frac{1}{2}\theta^{il}\eta_{00}\eta_{lk}\partial_{i}(J^{0}\pi^{k})+\frac{1}{2}\theta^{il}\partial_{i}(J^{k}F_{kl}).$ (3.23)
Combining Eqs. (3.22) and (3.23) with Eq. (3.20) gives
$\dot{\phi}_{2}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}+\frac{1}{2}\theta^{jl}\partial_{i}\!\bigl(\partial_{l}(J^{i}A_{j})\bigr)$
$\quad+\theta^{li}\partial_{i}\!\bigl(J^{k}\partial_{l}A_{k}\bigr)+\frac{1}{2}\theta^{il}\partial_{i}\!\bigl(J^{k}\partial_{k}A_{l}\bigr)$
$\quad+\frac{1}{2}\theta^{kl}\partial_{l}\!\left(A_{k}\,\partial_{0}J^{0}\right).$ (3.24)
If current conservation is not imposed a priori, the consistency condition
$\Phi_{3}^{\rm cand}(x):=\dot{\phi}_{2}(x)\approx 0$ (3.25)
defines the third-stage candidate,
$\Phi_{3}^{\rm cand}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}$ (3.26)
$\quad+\theta^{li}\partial_{i}\!\bigl(J^{k}\partial_{l}A_{k}\bigr)$
$\quad+\frac{1}{2}\theta^{jl}\partial_{i}\!\bigl(\partial_{l}(J^{i}A_{j})\bigr)$
$\quad+\frac{1}{2}\theta^{il}\partial_{i}\!\bigl(J^{k}\partial_{k}A_{l}\bigr)$
$\quad+\frac{1}{2}\theta^{kl}\partial_{l}\!\left(A_{k}\,\partial_{0}J^{0}\right).$
At $\theta=0$, this reduces to
$\Phi_{3}^{\rm cand}\big|_{\theta=0}=-\partial_{\mu}J^{\mu},$ (3.27)
so the tertiary-stage candidate collapses to the standard current-compatibility condition. The first-order deformation of this condition—and, as we now show, its Lagrangian meaning—is the central technical result of the paper.
3.3 Main identity: tertiary candidate as the canonical pullback of $\partial_{\nu}\mathcal{E}_{B}^{\nu}$
The goal of this subsection is to establish the main identity of the paper: the tertiary-stage candidate $\Phi_{3}^{\rm cand}=\dot{\phi}_{2}$ obtained above by preservation of the secondary Gauss-type constraint coincides, once expressed in the same $\mathcal{O}(\theta)$ canonical variables, with the divergence $\partial_{\nu}\mathcal{E}_{B}^{\nu}$ of the mapped Euler–Lagrange equations. This identity fixes the canonical locus at which the source-induced obstruction of the action-level Banerjee realization enters the Dirac chain.
Proposition 1.
To first order in $\theta$, and without imposing $\partial_{\mu}J^{\mu}=0$, the third-stage object generated by preserving the secondary Gauss-type constraint satisfies the exact algebraic identity
$\boxed{\;\partial_{\nu}\mathcal{E}_{B}^{\nu}=\dot{\phi}_{2}=\Phi_{3}^{\rm cand}\;}$ (3.28)
once the divergence of the mapped Euler–Lagrange equations is rewritten in the same canonical phase-space variables used in the Dirac analysis and the comparison is consistently truncated at first order in $\theta$. In the action-level Banerjee realization considered here, the tertiary-stage candidate is therefore algebraically identical to the canonical phase-space pullback of the divergence of the mapped Euler–Lagrange equations. This equality does not imply that the two objects have the same geometric status; it states only that, when expressed in the same $\mathcal{O}(\theta)$ canonical framework, they encode the same source-compatibility content.
Proof sketch. A direct Lagrangian computation is combined with an independent canonical computation. On the Lagrangian side, starting from the Banerjee-mapped Lagrangian in Eq. (3.1),
$\mathcal{L}_{B}=\mathcal{L}_{\text{ws}}(F)+\mathcal{L}_{J}(A,\partial A;J),$ (3.29)
the Euler–Lagrange operator decomposes as
$\mathcal{E}_{B}^{\nu}=\mathcal{E}_{\text{ws}}^{\nu}+\mathcal{E}_{J}^{\nu},$ (3.30)
with source contribution
$\mathcal{E}_{J}^{\nu}={}$ $-J^{\nu}-\theta^{\nu\alpha}J^{\beta}F_{\alpha\beta}-\theta^{\alpha\beta}(\partial_{\alpha}J^{\nu})A_{\beta}$ (3.31)
$\quad-\theta^{\alpha\beta}J^{\nu}\partial_{\alpha}A_{\beta}+\frac{1}{2}\theta^{\nu\beta}(\partial_{\mu}J^{\mu})A_{\beta}.$
Since $\mathcal{L}_{\text{ws}}$ depends on $A_{\mu}$ only through the antisymmetric field strength $F_{\mu\nu}$, its Euler–Lagrange operator has the form
$\mathcal{E}_{\text{ws}}^{\nu}=\partial_{\mu}P^{\mu\nu},\qquad P^{\mu\nu}:=\frac{\partial\mathcal{L}_{\text{ws}}}{\partial(\partial_{\mu}A_{\nu})},$ (3.32)
with $P^{\mu\nu}=-P^{\nu\mu}$. Therefore
$\partial_{\nu}\mathcal{E}_{\text{ws}}^{\nu}=0,$ (3.33)
and hence
$\partial_{\nu}\mathcal{E}_{B}^{\nu}=\partial_{\nu}\mathcal{E}_{J}^{\nu}.$ (3.34)
On the canonical side, the computation of $\dot{\phi}_{2}$ carried out in Sec. 3.2 yields Eq. (3.26). Using the antisymmetry of $\theta^{ij}$ to rearrange terms, both $\partial_{\nu}\mathcal{E}_{B}^{\nu}$ and $\dot{\phi}_{2}$ reduce to the common expression
$\Phi_{3}^{\rm cand}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}+\theta^{li}\partial_{i}(J^{k}\partial_{l}A_{k})$ (3.35)
$\quad-\frac{1}{2}\theta^{il}\partial_{i}\!\bigl[A_{l}(\partial_{\mu}J^{\mu})\bigr].$
A fully independent derivation—which does not rely on the compact form above and uses only Eqs. (3.19)–(3.20) together with the functional derivatives of $\phi_{2}$—is provided in Appendix C. $\square$
Equation (3.35) fixes the Lagrangian meaning of the third-stage candidate: within the action-level Banerjee realization, $\Phi_{3}^{\rm cand}$ is the canonical phase-space pullback of $\partial_{\nu}\mathcal{E}_{B}^{\nu}$, without implying any stronger equivalence between their geometric roles. An important structural remark is that this identity is map-dependent: the $F_{\alpha\beta}$-dependent terms specific to the Banerjee parametrization $(c_{1},c_{2},c_{3})=(1,\tfrac{1}{2},0)$ enter the explicit form of $\Phi_{3}^{\rm cand}$ through the source coupling $\mathcal{L}_{J}$. For a different SW current map—such as the minimal map of Adorno et al. [12]—the same Dirac–Bergmann procedure would produce a tertiary candidate with different $\mathcal{O}(\theta)$ coefficients. The identity (3.28) therefore characterizes the canonical fingerprint of the Banerjee realization specifically, rather than a universal feature of any action-level SW-mapped theory with external sources.
Equation (3.28) also permits a precise comparison with the canonical external-source analysis of Cabo and Shabad [13]: in their construction, the source-dependent condition selected by consistency appears at the secondary stage, whereas in the present framework the corresponding source-compatibility content is generated at the tertiary stage through $\Phi_{3}^{\rm cand}$. The comparison is therefore functional rather than stage-by-stage: the two objects are not being identified with one another, but in both analyses constraint preservation isolates the source-dependent condition required for consistency in the presence of external currents.
3.4 Preservation of $\Phi_{3}^{\rm cand}$ and generic closure by multiplier fixing
Having identified $\Phi_{3}^{\rm cand}$ as the next consistency condition, the Dirac algorithm requires
$\dot{\Phi}_{3}^{\rm cand}(x)={}$ $\{\Phi_{3}^{\rm cand}(x),H_{T}\}$ (3.36)
$\quad+\frac{\partial\Phi_{3}^{\rm cand}(x)}{\partial t}\Big|_{\rm expl}\approx 0.$
Two structural facts are immediate. Since $\Phi_{3}^{\rm cand}$ contains no canonical momenta,
$\{\Phi_{3}^{\rm cand}(x),\Phi_{3}^{\rm cand}(y)\}=0,$ (3.37)
and the only $A_{0}$ dependence of $\Phi_{3}^{\rm cand}$ is through the term $\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}$, so its bracket with the primary constraint is
$\{\Phi_{3}^{\rm cand}(x),\phi_{1}(y)\}$ $=\Theta^{l}(x)\,\partial_{l}^{x}\delta^{3}(x-y),$ (3.38)
$\Theta^{l}$ $=\theta^{li}\partial_{i}J^{0}.$
Therefore the generic structure of the next consistency equation is
$\dot{\Phi}_{3}^{\rm cand}={}$ $\Theta^{l}\partial_{l}u_{1}+\mathcal{F}[A_{\mu},\pi^{\mu};J^{\mu},\partial J,\partial^{2}J]$ (3.39)
$\quad-\partial_{0}(\partial_{\mu}J^{\mu})\approx 0,$
where $\mathcal{F}$ collects the remaining first-order terms. The explicit expression is recorded in Appendix B. Generically, for $\Theta^{l}\neq 0$, Eq. (3.39) fixes the primary multiplier $u_{1}$ along the differential operator $\Theta^{l}\partial_{l}$ rather than generating an independent quaternary constraint. In other words, the generic inhomogeneous branch closes by multiplier fixing at the next stage. Only in degenerate sectors, such as $\Theta^{l}=0$ together with additional cancellations, can the algorithm continue further.
3.5 Equal-time algebra and hierarchical source-kernel structure
For the unrestricted external-current problem, it is useful to record the equal-time algebra established up to the tertiary stage. Besides $\{\phi_{1},\phi_{1}\}=0$ and $\{\phi_{1},\phi_{2}\}=0$, one has
$\{\phi_{2}(x),\phi_{2}(y)\}={}$ $\,\frac{1}{2}\,\Theta^{k}(y)\,\partial_{k}^{y}\delta^{3}(x-y)$ (3.40)
$\quad-\frac{1}{2}\,\Theta^{k}(x)\,\partial_{k}^{x}\delta^{3}(x-y),$
while Eq. (3.38) gives, by antisymmetry,
$\{\phi_{1}(x),\Phi_{3}^{\rm cand}(y)\}=-\Theta^{l}(y)\,\partial_{l}^{y}\delta^{3}(x-y),$ (3.41)
and
$\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}{}$ $=\Sigma^{lk}(y)\,\partial_{l}^{y}\partial_{k}^{y}\delta^{3}(x-y)$ (3.42)
$+\Xi^{l}(y)\,\partial_{l}^{y}\delta^{3}(x-y),$
with
$\Sigma^{lk}$ $=\frac{1}{2}\Bigl(\theta^{li}\partial_{i}J^{k}+\theta^{ki}\partial_{i}J^{l}\Bigr),$ (3.43)
$\Xi^{l}$ $=-\frac{1}{2}\,\theta^{il}\,\partial_{i}(\partial_{\mu}J^{\mu}),$
together with
$\{\Phi_{3}^{\rm cand}(x),\Phi_{3}^{\rm cand}(y)\}=0.$ (3.44)
Equations (3.40)–(3.43) reveal that the source dependence of the tertiary stage is not encoded in a single obstruction, but in three distinct kernels. The kernel $\Theta^{l}$ controls both the secondary self-bracket and the primary–tertiary bracket, whereas the mixed secondary–tertiary bracket depends on $\Sigma^{lk}$ and $\Xi^{l}$. These kernels probe different aspects of the prescribed source profile: $\Theta^{l}$ measures the variation of the charge density along the non-commutative directions, $\Sigma^{lk}$ measures the symmetrized spatial gradients of the current flow, and $\Xi^{l}$ measures the spatial gradient of the source divergence. They are not redundant: ordinary current conservation forces $\Xi^{l}=0$ but does not by itself imply either $\Theta^{l}=0$ or $\Sigma^{lk}=0$; likewise, spatial constancy of $J^{k}$ may set $\Sigma^{lk}=0$ without constraining the remaining kernels.
This observation induces a hierarchical organization of source sectors. The generic non-degenerate unrestricted branch is characterized by $\Theta^{l}\neq 0$. In that case, the source-induced obstruction is already visible in $\{\phi_{2},\phi_{2}\}$ and $\{\phi_{1},\Phi_{3}^{\rm cand}\}$, and preservation of $\Phi_{3}^{\rm cand}$ constrains the primary multiplier through Eq. (3.39); subject to the properties of the differential operator $\Theta^{l}\partial_{l}$, the algorithm closes by multiplier determination rather than by an independent quaternary constraint. No unrestricted first-class interpretation is supported in this sector. By contrast, when $\Theta^{l}=0$ but at least one of $\Sigma^{lk}$ or $\Xi^{l}$ remains non-vanishing, the obstruction disappears from $\{\phi_{2},\phi_{2}\}$ and $\{\phi_{1},\Phi_{3}^{\rm cand}\}$ but survives in the mixed bracket $\{\phi_{2},\Phi_{3}^{\rm cand}\}$, so the vanishing of $\Theta^{l}$ alone is not sufficient to recover a first-class structure.
At the stage reached here, the manuscript does not claim a complete first-/second-class classification of the unrestricted source sector. What has been established is the equal-time algebra available up to the tertiary stage and the associated consistency structure, which should be read as the canonical information presently available for the unrestricted problem rather than as an exhaustive classification of the external-current theory.
Regime change and outlook for the reduced description. The analysis of Secs. 3.43.5 produces two structurally distinct outcomes. For the generic non-degenerate inhomogeneous branch $\Theta^{l}\neq 0$, preservation of $\Phi_{3}^{\rm cand}$ determines the primary multiplier along $\Theta^{l}\partial_{l}$ rather than yielding an independent quaternary constraint, and no standard first-class classification is available without further input. For source profiles satisfying simultaneously
$\Theta^{l}=0,\qquad\Sigma^{lk}=0,\qquad\Xi^{l}=0,\qquad\Phi_{3}^{\rm cand}\equiv 0,$ (3.45)
the displayed equal-time obstructions disappear, the surviving constraint pair $(\phi_{1},\phi_{2})$ becomes first-class in the strict equal-time sense, and the standard machinery of gauge generators, Dirac brackets, and reduced dynamics applies unchanged.
To see what these conditions select, consider the concrete case of purely $x^{1}$$x^{2}$ non-commutativity, $\theta^{12}=-\theta^{21}=\theta$ with all other components zero. The kernel $\Theta^{l}=\theta^{li}\partial_{i}J^{0}$ then requires $\partial_{1}J^{0}=\partial_{2}J^{0}=0$: the charge density must be uniform across the non-commutative plane. Likewise, $\Sigma^{lk}=0$ forces all spatial current components to be independent of $x^{1}$ and $x^{2}$, and $\Xi^{l}=0$ demands the same of $\partial_{\mu}J^{\mu}$. Any source of the form
$J^{\mu}=\bigl(\rho(t,x^{3}),\,0,\,0,\,j(t,x^{3})\bigr),\qquad\partial_{0}\rho+\partial_{3}j=0,$ (3.46)
automatically satisfies the first three conditions in Eq. (3.45). Under these conditions, the remaining requirement $\Phi_{3}^{\rm cand}\equiv 0$ reduces to ordinary current conservation $\partial_{\mu}J^{\mu}=0$, which is precisely the continuity equation already imposed in Eq. (3.46). The physical picture is therefore a conserved charge–current system whose profile may vary along the direction parallel to $\boldsymbol{\theta}$ but is uniform across the non-commutative plane. In particular, the sector conditions do not require $J^{\mu}$ to vanish: they restrict its derivative profile (spatial gradients and divergence) in the non-commutative directions, while the pointwise values $J^{0}$, $J^{k}$ remain non-trivial. This distinction is important for reading the reduced Hamilton equations of Sec. 4.3, where $J^{\mu}$ appears in its general functional form even though the brackets through which the equations are derived exist only for source profiles of the type (3.46).
The remainder of the paper, Sec. 4, develops this reduced description in that restricted subcase and uses it to confront the Banerjee route with the gauge-invariant equation-level formulation of Ref. [12]. No claim about the unrestricted external-current theory is made in what follows; the sectorial caveat is declared once here and not repeated subsection by subsection.
4 Reduced phase-space description in the restricted first-class sector
All constructions in this section presuppose the sector conditions of Eq. (3.45), under which the surviving equal-time constraint pair is $(\phi_{1},\phi_{2})$ and no further independent consistency condition arises at the level analyzed in Sec. 3. The reduced Hamiltonian structure, gauge generator, Dirac brackets and reduced dynamics obtained below are therefore claimed only for this restricted sufficient subcase and are not presented as statements about the unrestricted external-current problem.
4.1 Sector conditions, gauge generator, and degree-of-freedom count
Within the restricted sector defined by Eq. (3.45), the gauge generator associated with the surviving first-class pair $(\phi_{1},\phi_{2})$ can be introduced in the standard Castellani-type form
$G[\epsilon_{0},\epsilon]=\int d^{3}x\,\bigl[\epsilon_{0}(x)\phi_{1}(x)+\epsilon(x)\phi_{2}(x)\bigr],$ (4.1)
where $\epsilon_{0}$ and $\epsilon$ are arbitrary test functions at fixed time. Using the canonical equal-time brackets, the gauge variations of the basic variables read
$\delta A_{0}$ $=\epsilon_{0},$ $\delta A_{i}$ $=\partial_{i}\epsilon,$
$\delta\pi^{0}$ $=0,$ $\delta\pi^{i}$ $=\frac{1}{2}\theta^{ik}\partial_{k}(\epsilon J^{0}),$ (4.2)
and reduce to the standard canonical $U(1)$ transformations in the commutative limit. These transformations describe the residual equal-time gauge freedom of the dynamical sector in the restricted setting.
The independent physical variables are the transverse components of the gauge field and their conjugate momenta. Starting from the eight canonical field variables $(A_{\mu},\pi^{\mu})$, the surviving pair $(\phi_{1},\phi_{2})$ provides two first-class constraints, so that the reduced number of physical phase-space degrees of freedom is
$N_{\mathrm{phys}}=\frac{1}{2}(8-2\times 2)=2,$ (4.3)
corresponding to two propagating configuration-space degrees of freedom. The explicit gauge fixing, the corresponding Dirac reduction, and the resulting reduced brackets are addressed in Sec. 4.2.
4.2 Dirac brackets in Coulomb gauge
To perform the reduction, we impose the Coulomb-type gauge-fixing conditions
$\chi_{1}:=A_{0}\approx 0,\qquad\chi_{2}:=\partial_{i}A^{i}\approx 0.$ (4.4)
Together with the surviving pair $(\phi_{1},\phi_{2})$, these conditions define the second-class set
$\Gamma_{a}=(\phi_{1},\phi_{2},\chi_{1},\chi_{2}),$ (4.5)
with
$\phi_{1}=\pi^{0},\qquad\phi_{2}=\partial_{i}\pi^{i}-J^{0}+\frac{1}{2}\theta^{kl}\partial_{l}(A_{k}J^{0}).$ (4.6)
In the restricted sector, the non-vanishing equal-time brackets among these quantities are the standard ones,
$\{\phi_{1}(x),\chi_{1}(y)\}$ $=-\delta^{3}(x-y),$ (4.7)
$\{\phi_{2}(x),\chi_{2}(y)\}$ $=-\nabla^{2}\delta^{3}(x-y),$
while the source-dependent obstructions displayed in the unrestricted problem have already vanished. The corresponding constraint matrix is therefore invertible in the standard equal-time sense.
The associated Dirac bracket is
$\{F,G\}_{D}={}$ $\{F,G\}-\int d^{3}u\,d^{3}v\,\{F,\Gamma_{a}(u)\}$
$\qquad\qquad\,C^{-1}_{ab}(u,v)\,\{\Gamma_{b}(v),G\},$
where $C_{ab}(x,y)=\{\Gamma_{a}(x),\Gamma_{b}(y)\}$ and $C^{-1}_{ab}$ denotes its inverse. The resulting brackets among the independent transverse variables take the familiar Coulomb-gauge form,
$\{A_{i}(x),A_{j}(y)\}_{D}=0,$ (4.8)
$\{\pi_{i}(x),\pi_{j}(y)\}_{D}=0,$ (4.9)
and
$\{A_{i}(x),\pi_{j}(y)\}_{D}=\left(\delta_{ij}-\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\right)\delta^{3}(x-y).$ (4.10)
Accordingly, only the transverse components of the gauge field and canonical momentum remain as independent reduced variables. These brackets provide the canonical input for the reduced Hamiltonian evolution discussed in the next subsection.
4.3 Reduced Hamilton equations and Maxwell-like rewriting
With the Dirac brackets in place, the reduced Hamiltonian generates the equal-time evolution of the independent phase-space variables through
$\dot{A}_{s}$ $=\{A_{s},H_{c}\}_{*}\,$, (4.11)
$\dot{\pi}^{s}$ $=\{\pi^{s},H_{c}\}_{*}\,.$ (4.12)
A remark on the logical status of the source terms is in order. The Dirac bracket $\{,\}_{*}$ constructed in Sec. 4.2 exists only in the restricted sector (3.45), because the invertibility of the constraint matrix in its standard Coulomb-gauge form requires $\Theta^{l}=\Sigma^{lk}=\Xi^{l}=0$. The canonical Hamiltonian $H_{c}=H_{\text{ws}}+H_{J}$, however, retains the full source coupling: the sector conditions constrain the admissible profiles of $J^{\mu}$—its spatial derivatives and divergence—but do not set $J^{\mu}$ itself to zero. The resulting Hamilton equations therefore contain the external current in its general functional form, while remaining valid only for source profiles satisfying Eq. (3.45). For concreteness, the admissible class includes the conserved, NC-plane-uniform profiles illustrated in Eq. (3.46): the source terms $J^{0}$, $J^{k}$ in the equations below are non-vanishing but have restricted spatial gradients. The equations are presented in this unsimplified form to make the comparison with the Euler–Lagrange system of Sec. 4.4 as direct as possible; no further reduction using the sector conditions on the source terms is attempted here.
At first order in $\theta$, the reduced evolution is most compactly written as
$\dot{A}_{s}=\eta^{sc}\,\Lambda_{c}\,$, (4.13)
with
$\Lambda^{c}={}$ $-\eta_{00}\Bigl[\bigl(1-\mathbf{B}\cdot\boldsymbol{\theta}\bigr)\pi^{c}+(\boldsymbol{\pi}\cdot\boldsymbol{\theta})B^{c}+(\mathbf{B}\cdot\boldsymbol{\pi})\theta^{c}\Bigr]$
$+\tfrac{1}{2}\eta_{00}J^{0}(\mathbf{A}\times\boldsymbol{\theta})^{c}\,.$ (4.14)
For the momentum one finds
$\dot{\pi}^{s}$ $=-\varepsilon^{nsa}\partial_{n}\Bigl[\bigl(1+\mathbf{B}\!\cdot\!\boldsymbol{\theta}\bigr)F^{ns}+(\boldsymbol{\pi}\!\cdot\!\boldsymbol{\theta})\pi_{a}$
$\qquad\quad\;-\tfrac{1}{2}\bigl(\eta_{00}\pi_{j}\pi^{j}-B_{j}B^{j}\bigr)\theta_{a}\Bigr]$
$\quad+J^{0}\theta^{sl}\pi_{l}-J^{j}\theta^{sl}\partial_{j}A_{l}-J^{s}\,.$ (4.15)
Equations (4.13)–(4.15) already show that the reduced electric sector is not obtained from the Maxwell relation $E^{i}=\pi^{i}$ alone, but from a source- and field-dependent deformation of that identification at first order in $\theta$. In the commutative limit, Eqs. (4.14) and (4.15) reduce to the standard Hamilton equations of Maxwell theory with external sources.
Two structural features of the source sector should be noted at the outset. First, the source terms in Eqs. (4.14) and (4.15) contain the gauge potential $A_{j}$ explicitly—not only through the field strength $F_{jl}$. In $\Lambda^{c}$, the term $\tfrac{1}{2}\eta_{00}J^{0}(\mathbf{A}\times\boldsymbol{\theta})^{c}$ depends on $A_{j}$ itself; in $\dot{\pi}^{s}$, the combination $J^{j}\theta^{sl}\partial_{j}A_{l}$ involves $\partial_{j}A_{l}$ rather than $F_{jl}=\partial_{j}A_{l}-\partial_{l}A_{j}$. These terms are the Hamiltonian counterpart of the gauge-non-covariant source coupling identified at the Lagrangian level by Adorno et al. [12]. Within the present gauge-fixed formulation ($A_{0}=0$, $\partial_{k}A^{k}=0$), the $A_{j}$ that appears is the physical transverse potential and the equations are self-consistent; but the appearance of the undifferentiated potential—rather than the field strength—in the source terms signals that the underlying action-level realization is not gauge covariant when external currents are present. The precise decomposition of this non-covariance into a scheme-dependent gauge-invariant piece and an explicitly gauge-dependent piece is carried out in the on-shell comparison of Sec. 4.5. Second, the field-strength bilinear sector (all terms not involving $J^{\mu}$) is manifestly gauge invariant, as it depends only on $F_{ij}$, $B^{i}$, and $\pi^{i}$. The asymmetry between gauge-invariant kinematics and gauge-non-covariant source coupling will be reflected in the on-shell comparison, where the Gauss sector is recovered exactly while the Ampère sector retains irreducible residues.
Maxwell-like rewriting of the reduced dynamics. The reduced Hamilton equations admit a Maxwell-like rewriting that is convenient for the comparison with the mapped Euler–Lagrange system carried out in Secs. 4.44.5. Using Eq. (2.2) and the solution (3.11) for $F^{0i}$, the electric field may be written as
$E^{i}={}$ $\,\pi^{k}F_{kl}\theta^{li}+\pi^{k}F_{ml}\eta^{im}\eta_{nk}\theta^{ln}$ (4.16)
$\quad-\pi^{i}\left(\tfrac{1}{2}F_{kl}\theta^{lk}-1\right)+\tfrac{1}{2}\,\theta^{ji}A_{j}J^{0}\,.$
Note that the last term, $\tfrac{1}{2}\theta^{ji}A_{j}J^{0}$, makes the effective electric field explicitly dependent on the gauge potential in the presence of sources; in the source-free limit $J^{0}\to 0$ this term vanishes and $E^{i}$ becomes a function of $F_{\mu\nu}$ and $\pi^{k}$ alone. Hence
$\dot{A}_{s}=\eta_{00}E_{s}\,.$ (4.17)
The magnetic equations remain kinematic:
$\partial_{0}\mathbf{B}=-\nabla\times\mathbf{E}\,$, (4.18)
and
$\nabla\cdot\mathbf{B}=0\,.$ (4.19)
The physical Gauss law should not be inferred from the divergence of Eq. (4.17). That equation governs the time evolution of the spatial gauge potential $A_{i}$, whereas Gauss’s law belongs to the canonical constraint structure of the theory. In the present Hamiltonian formulation, the Gauss condition is encoded directly in the secondary constraint obtained from preserving the primary constraint $\pi^{0}\approx 0$, namely
$\partial_{i}\pi^{i}=J^{0}-\frac{1}{2}\,\theta^{kl}\,\partial_{l}(A_{k}J^{0})\,.$ (4.20)
Thus, Gauss’s law is not an independent consequence of the evolution equation for $A_{i}$, but a restriction that defines the admissible phase-space configurations at each time slice. Its role is to constrain the longitudinal part of the canonical momentum in the presence of the prescribed external charge density, including the first-order non-commutative correction.
Identifying the reduced electric-displacement variable with the canonical momentum,
$D^{s}\equiv\pi^{s},$ (4.21)
Eq. (4.15) can be rewritten in Maxwell-like form as
$(\nabla\times\mathbf{H})^{s}=T^{s}-\partial_{0}D^{s},$ (4.22)
with
$\mathbf{H}$ $=\bigl[1+B_{j}\theta^{j}\bigr]\mathbf{B}+\eta_{00}(\boldsymbol{\pi}\cdot\boldsymbol{\theta})\boldsymbol{\pi}$
$\quad+\tfrac{1}{2}(B_{a}B^{a})\boldsymbol{\theta}-\tfrac{1}{2}\eta_{00}(E_{j}E^{j})\boldsymbol{\theta}\,$, (4.23)
$T^{s}$ $=J^{0}\,\theta^{sl}\,\pi_{l}-J^{j}\,\theta^{sl}\,\partial_{j}A_{l}-J^{s}\,.$ (4.24)
Equation (4.22) is not an independent constitutive law but a compact rewriting of the reduced Hamilton equation for $\pi^{s}$; the notation is organizational. The source sector $T^{s}$ contains the external current in its general functional form; the sector conditions (3.45) constrain only the derivative profile of $J^{\mu}$, not its pointwise values, so $J^{0}$, $J^{j}$, $J^{s}$ are not simplified here. Even so, $T^{s}$ already contains both gauge-invariant and explicitly gauge-dependent contributions, which anticipates the asymmetric on-shell comparison with the Euler–Lagrange equations: the Gauss sector is anchored directly in the secondary constraint, whereas the Ampère-like sector inherits the full source-dependent deformation of the reduced Hamiltonian evolution. The term $\eta_{00}(\boldsymbol{\pi}\cdot\boldsymbol{\theta})\boldsymbol{\pi}$ in $\mathbf{H}$ encodes the anisotropic deformation of the effective magnetic permeability induced by the NC structure: with $\eta_{00}=+1$ and $\boldsymbol{\pi}\to\mathbf{E}$ on shell, this becomes $(\mathbf{E}\cdot\boldsymbol{\theta})\mathbf{E}$, coupling the electric sector to the NC direction $\boldsymbol{\theta}$ already at the constitutive level. The commutative limit $\theta\to 0$ recovers $\mathbf{H}=\mathbf{B}$, as required.
This Maxwell-like representation should not be overinterpreted. In particular, the source sector $T^{s}$ is not gauge covariant: it contains the gauge potential $A_{j}$ through the term $J^{j}\theta^{sl}\partial_{j}A_{l}$, which does not combine into $F_{jl}$. Accordingly, the Ampère-like equation (4.22) does not have the same gauge-invariance status as its commutative counterpart. The representation is nevertheless useful because it separates the gauge-invariant field-strength block ($\mathbf{H}$, $D^{s}$) from the gauge-non-covariant source block ($T^{s}$) in a form that makes the comparison with the mapped Euler–Lagrange structure (Sec. 4.4) and the on-shell gauge-covariance analysis (Sec. 4.5) as direct as possible. Its meaning is therefore internal to the restricted sufficient first-class subcase.
4.4 Comparison with the Banerjee Euler–Lagrange system (off-shell)
Having rewritten the reduced Hamilton equations in a Maxwell-like form, we now compare them with the Euler–Lagrange equations of the mapped theory. The point of the comparison is limited but structural: it tests how far the reduced canonical dynamics reproduces the field-equation content of the mapped Banerjee theory once the source-dependent obstructions have been removed by the sector conditions of Eq. (3.45). The logical relation among the relevant formulations is
[…]
Here the solid arrows denote exact equivalence or direct comparison inside the restricted branch, whereas the dashed arrow denotes only a partial on-shell correspondence between the action-level Banerjee route and the gauge-covariant equation-level route of Adorno et al.
Applying the Euler–Lagrange operator to $\mathcal{L}=\mathcal{L}_{\text{ws}}+\mathcal{L}_{J}$, one obtains
$\mathcal{E}^{\nu}_{B}$ $=\partial_{\mu}F^{\mu\nu}+\theta^{\alpha\beta}\Bigl[\tfrac{1}{4}\partial^{\nu}(F_{\alpha\mu}F_{\beta}{}^{\mu})-\partial_{\mu}(F^{\mu}_{\ \alpha}F^{\nu}_{\ \beta})$
$+F_{\mu\alpha}\partial_{\beta}F^{\mu\nu}\Bigr]-J^{\nu}-\theta^{\nu\alpha}J^{\beta}F_{\alpha\beta}-\theta^{\alpha\beta}(\partial_{\alpha}J^{\nu})A_{\beta}$
$-\theta^{\alpha\beta}J^{\nu}\partial_{\alpha}A_{\beta}+\tfrac{1}{2}\theta^{\nu\beta}(\partial_{\mu}J^{\mu})A_{\beta}=0\,.$ (4.26)
The first two lines constitute the common field-strength bilinear sector (same in both the Banerjee and Adorno formulations); the remaining source-dependent terms contain explicit gauge potentials and are therefore gauge-non-covariant for external fixed currents. The term $\tfrac{1}{2}\theta^{\nu\beta}(\partial_{\mu}J^{\mu})A_{\beta}$ vanishes when the external current is conserved but must be retained in the general off-shell expression.
For comparison, Adorno et al. obtain, with the minimal current map,
$\mathcal{E}^{\nu}_{A}$ $=\partial_{\mu}\Bigl[f^{\mu\nu}\Bigl(1+\tfrac{g}{2}\theta^{\alpha\beta}f_{\alpha\beta}\Bigr)\Bigr]$
$\quad-g\theta^{\alpha\beta}\partial_{\mu}(f^{\mu\nu}f_{\alpha\beta})-g\theta^{\alpha\beta}\partial_{\mu}(f^{\mu\alpha}f^{\nu\beta})$
$\quad+g\theta^{\alpha\beta}\partial_{\beta}(f^{\mu\nu}f_{\mu\alpha})-\tfrac{g}{4}\theta^{\alpha\beta}\partial_{\alpha}(f_{\mu\nu}f^{\mu\nu})$
$\quad-\frac{4\pi}{c}\,j^{\nu}\Bigl(1+\tfrac{g}{2}\theta^{\alpha\beta}f_{\alpha\beta}\Bigr)$
$\quad-\frac{4\pi g}{c}\theta^{\alpha\beta}\Bigl[f_{\alpha\beta}j^{\nu}+f^{\nu}{}_{\alpha}j_{\beta}$
$\qquad\quad+A_{\alpha}\partial_{\beta}j^{\nu}-\tfrac{1}{2}A^{\nu}\partial_{\alpha}j_{\beta}\Bigr]=0\,.$ (4.27)
The field-strength sector agrees in both formulations because it comes from the same SW-mapped Maxwell block $\mathcal{L}_{\text{ws}}$. The difference lies entirely in the source sector: the Banerjee map generates, in addition to $J^{\nu}$, the gauge-non-covariant terms $\theta^{\nu\alpha}J^{\beta}F_{\alpha\beta}$, $\theta^{\alpha\beta}(\partial_{\alpha}J^{\nu})A_{\beta}$, and $\theta^{\alpha\beta}J^{\nu}\partial_{\alpha}A_{\beta}$. The last two contain the gauge potential explicitly and produce the gauge-covariance obstruction identified by Adorno et al. Thus, the comparison with Adorno must be understood as a comparison between different admissible SW-map choices, not as an identity between the two Lagrangian formulations. In the present restricted sector, the reduced Hamilton equations reproduce the common field-strength content of the mapped Euler–Lagrange system, while the residual source-dependent mismatch measures the limitation of the reduced Banerjee description rather than an inconsistency of the unrestricted theory as a whole.
4.5 On-shell comparison in the reduced regime
Section 4.4 compared the reduced Hamilton equations with the mapped Euler–Lagrange equations at the level of their explicit component form. We now evaluate that comparison on shell, under the gauge conditions $A_{0}=0$ and $\partial_{k}A^{k}=0$, working throughout to first order in $\theta$. The temporal and spatial sectors do not behave symmetrically: the temporal comparison is anchored directly in the secondary constraint, whereas the spatial comparison inherits the full source-dependent deformation of the reduced Hamiltonian evolution.
Spatial component: partial correspondence. From Eq. (3.8), the canonical momentum on shell is
$\pi^{i}=F^{0i}\bigl[1+\mathcal{O}(\theta)\bigr]+\tfrac{1}{2}\,\theta^{ji}A_{j}J^{0}\,.$ (4.28)
Substituting into Eq. (4.15), the field-strength sector yields exactly the bilinear structure appearing on the left-hand side of the gauge-invariant field equation of Ref. [12]. This part of the correspondence is exact within the restricted sufficient first-class regime.
The source contribution is
$T^{s}=\,J^{0}\theta^{sl}\pi_{l}-J^{j}\theta^{sl}\partial_{j}A_{l}-J^{s}\,$, (4.29)
which can be rewritten as
$T^{s}=\,\bigl(J^{\mu}\theta^{sl}F_{\mu l}-J^{s}\bigr)-J^{j}\theta^{sl}\partial_{l}A_{j}\,.$ (4.30)
Two independent obstructions remain:
(i) Scheme-dependent gauge-invariant obstruction. The term $J^{\mu}\theta^{sl}F_{\mu l}$ is gauge invariant but does not appear on the right-hand side of the gauge-invariant equation-level formulation of Ref. [12]. It reflects the Banerjee-map choice and cannot be removed by the constraint projection.
(ii) Gauge-dependent obstruction. The term $J^{j}\theta^{sl}\partial_{l}A_{j}$ contains $A_{j}$ explicitly and remains gauge dependent. Under $A_{j}\to A_{j}+\partial_{j}\Lambda$ it transforms as
$-J^{j}\theta^{sl}\partial_{l}A_{j}$ $\;\longrightarrow\;-J^{j}\theta^{sl}\partial_{l}A_{j}-J^{j}\theta^{sl}\partial_{l}\partial_{j}\Lambda\,$, (4.31)
and the residual term does not vanish generically. Hence the spatial Hamiltonian equation does not reduce to the gauge-invariant Ampère equation of Ref. [12]: what survives is an exact identification of the common field-strength bilinear sector, but not a full reduction of the spatial Hamiltonian equation to the gauge-invariant Ampère equation.
Temporal component: exact Gauss-law recovery. The temporal component is different. On the constraint surface,
$\partial_{i}\pi^{i}=J^{0}-\tfrac{1}{2}\theta^{kl}\partial_{l}(A_{k}J^{0})\,.$ (4.32)
Using the decomposition (3.9)–(3.10) of the spatial canonical momentum introduced in Sec. 3.1, one obtains
$\begin{split}\partial_{i}\pi^{i}={}&\partial_{i}F^{0i}+\partial_{i}\mathcal{K}^{i}(F)+\tfrac{1}{2}\theta^{ji}(\partial_{i}A_{j})\,J^{0}\\ &+\tfrac{1}{2}\theta^{ji}A_{j}\,\partial_{i}J^{0}\,.\end{split}$ (4.33)
The two source-dependent terms in Eq. (4.33) cancel exactly against the corresponding terms on the right-hand side of Eq. (4.32): the $(\partial_{i}A_{j})J^{0}$ pair cancels by antisymmetry of $\theta^{ji}$, and the $A_{k}\partial_{l}J^{0}$ pair cancels identically. The remainder contains only the field-strength bilinear block $\partial_{i}F^{0i}+\partial_{i}\mathcal{K}^{i}(F)$ and no explicit potential terms, and therefore coincides with the $\mu=0$ component of the gauge-invariant formulation of Ref. [12]. This is the exact on-shell result within the restricted sufficient first-class regime.
NC-deformed conservation on the constraint surface.
By Proposition 1, requiring the third-stage consistency condition $\Phi_{3}^{\rm cand}\approx 0$ is equivalent to requiring that the mapped field equations be divergence-consistent once their divergence is rewritten in the same canonical phase-space variables and truncated consistently at first order in $\theta$. At first order this amounts to a source-compatibility requirement generated on shell by the algorithm, not to ordinary current conservation imposed from outside. In the restricted sufficient first-class subcase studied here, no separate reduced tertiary representative is introduced: one must instead impose directly that the full tertiary condition vanish identically together with the kernel conditions displayed in Sec. 3.5.
Taken together, the main-text identity of Sec. 3.3 and the on-shell comparison above locate the source-induced obstruction of the action-level Banerjee realization in two complementary places: at the tertiary stage of the Dirac chain for the unrestricted problem, and in the spatial Ampère-type sector of the reduced dynamics for the restricted first-class subcase. The temporal sector, anchored directly in the secondary constraint, reproduces the gauge-invariant Gauss law exactly on shell; the spatial sector retains two irreducible residues—one scheme-dependent and gauge invariant, inherited from the current-map ambiguity, and one explicitly gauge dependent. The following section examines what this canonical diagnosis implies for the Lagrangian mismatch identified by Adorno et al., and delimits the scope of the present results with respect to alternative SW current maps and action-level realizations.
5 Discussion
5.1 Canonical versus Lagrangian locus of the obstruction
The analysis of Sec. 3 assigns a precise canonical meaning to the source-induced obstruction. In the unrestricted problem, the preservation of the secondary Gauss-type constraint produces the third-stage identity (3.28). Its zeroth-order part is the standard compatibility term $\partial_{\mu}J^{\mu}$, while its full first-order content is the NC-deformed source-compatibility condition generated on shell by the algorithm. The crucial point is that this quantity is simultaneously the next object in the Dirac chain and, after canonical pullback, the divergence of the Banerjee Euler–Lagrange equations. The generic source sector therefore closes by multiplier fixing rather than by the appearance of a universal quaternary constraint, while first-class reduced behavior survives only as a derived special subcase. The Hamiltonian analysis thus sharpens the Lagrangian obstruction identified by Adorno et al.: what appears there as a failure of gauge covariance in the action-level route reappears here as the explicit point at which source compatibility enters the Dirac consistency chain. More precisely, the canonical formulation provides three layers of information that are not accessible from the Lagrangian side alone: it localizes the obstruction at a definite stage of the Dirac chain (the tertiary level, via the identity (3.28)); it classifies the source dependence through the hierarchical kernel structure $(\Theta^{l},\Sigma^{lk},\Xi^{l})$ of Sec. 3.5; and it decomposes the on-shell comparison into a Gauss sector that is recovered exactly and an Ampère sector that retains irreducible residues (Sec. 4.5). None of these distinctions has a direct counterpart in the equation-level analysis of Ref. [12].
The asymmetry between the temporal and spatial on-shell comparisons in Sec. 4.5 is not incidental. The temporal component is controlled by the secondary constraint itself, which by construction imposes the NC-deformed Gauss law. The spatial comparison, by contrast, requires the full reduced Hamilton equations for $\dot{\pi}^{s}$, whose source sector inherits both the Banerjee-map choice and the canonical deformation of $\pi^{i}$. Within the Banerjee realization studied here, it is precisely that map dependence—in particular the $F_{\alpha\beta}$-dependent terms absent in the minimal Adorno map—that produces the residual gauge-invariant obstruction in the Ampère sector. The present analysis therefore supports a limited conclusion: for this specific first-order action-level current map, the reduced spatial sector does not reproduce the gauge-covariant target equations of Ref. [12] for generic source profiles. No broader no-go statement is claimed for other SW current maps or for other action-level realizations not analyzed here.
5.2 Comparison with prior external-source analyses
A related structural remark concerns the role of a temporal source component. The Cabo–Shabad construction is formulated under the restriction $J_{4}^{\,a}=0$, which prevents the temporal source density from shifting Gauss’s law at the stage where their source-dependent secondary condition emerges. Had such a component been retained, the corresponding consistency equation would no longer remain purely spatial. The present SW-mapped model shows explicitly how that source-compatibility content can instead reappear at a later stage of the chain, through the tertiary-stage candidate $\Phi_{3}^{\rm cand}$.
The source-dependent comparison with earlier literature should therefore be read with some care. Cabo–Shabad remain the closest canonical analogue, because in both analyses a prescribed external current enters the consistency chain itself; the difference is that their non-Abelian construction is written with $J_{4}^{\,a}=0$, so the source-dependent condition appears already at the secondary stage, whereas here the temporal density $J^{0}$ and the spatial current $J^{k}$ survive together and the corresponding compatibility content is first isolated at the tertiary stage [13]. Przeszowski’s analysis of Yang–Mills theory with arbitrary external sources points in the same methodological direction: there too the source sector cannot be treated as canonically innocuous, and a modified source coupling is introduced before the quantization procedure is organized [14]. Sikivie and Weiss provide a different kind of comparison, since their static external sources are used to define physically distinct classical solution sectors rather than to study the first-/second-class constraint split [15]. By contrast, Kruglov’s NC Maxwell model employs a prescribed Abelian current at the Lagrangian level, but leaves aside the Dirac–Bergmann question of how the source deforms the canonical chain [16]. The gain of the present Hamiltonian analysis is precisely to make that source dependence explicit in a setting where the current is fixed, Abelian, SW-mapped, and allowed to carry both temporal and spatial components.
5.3 Scope, limitations, and outlook
The results reported here are confined to the first-order action-level Banerjee realization of NC $U(1)$ electrodynamics with purely space–space non-commutativity ($\theta^{0i}=0$) and a prescribed, non-dynamical external current $J^{\mu}$. Within that scope, the main identity (3.28) is algebraic: it holds as a statement in the $\mathcal{O}(\theta)$ canonical variables, without additional assumptions on the source profile. The generic closure by multiplier fixing (Sec. 3.4) is itself a structural result: it establishes that, for the Banerjee realization with $\Theta^{l}\neq 0$, the Dirac algorithm terminates without producing an independent quaternary constraint, and that the obstruction is absorbed into the primary multiplier along the differential operator $\Theta^{l}\partial_{l}$. This is not a limitation of the analysis but a positive characterization of how the unrestricted external-current theory closes canonically. The sectorial classification of Sec. 3.5 and the reduced phase-space construction of Sec. 4 are, by contrast, restricted: the gauge-generator, Dirac-bracket and Hamilton-equations sector-level statements rely on the sufficient conditions (3.45), and should not be transferred to generic source profiles without further analysis. Higher-order extensions in $\theta$, temporal non-commutativity $\theta^{0i}\neq 0$, and dynamical (rather than prescribed) sources each raise distinct conceptual issues that are not addressed here.
Because the central identity is map-dependent (see the remark following Proposition 1), two companion analyses are currently under investigation. The first extends the present Hamiltonian construction to the minimal current map of Ref. [12], in order to determine whether the residual gauge-invariant obstruction in the Ampère sector is specific to the Banerjee $F_{\alpha\beta}$-dependent terms or whether some weaker form of it survives in any action-level SW current map. The second constructs the canonical counterpart of the equation-level route itself, i.e. asks how the gauge-covariant Adorno field equations fit in a Dirac–Bergmann analysis when the SW map is applied after variation; a comparison between the two canonical constructions would make the relation between the two routes quantitative at the phase-space level. The approach of Refs. [21, 22] to SW-map ambiguities through local cohomology, together with the current-algebra perspective of Ref. [23], suggests natural organizing principles for such extensions. Neither task is required for the present claim, but together they would delimit how much of the obstruction identified here is tied to the Banerjee realization and how much is structural.
6 Conclusions
We studied the first-order Banerjee action-level realization of NC Maxwell theory with a fixed external current in the full Dirac–Bergmann phase space. The main result is that preservation of the Gauss-type secondary constraint produces a third-stage source-dependent object whose canonical form is algebraically identical, after canonical pullback, to the divergence of the mapped Euler–Lagrange equations. This identity is tied to the specific $F_{\alpha\beta}$-dependent terms of the Banerjee current map: it characterizes the canonical fingerprint of that particular realization, not a universal feature of any action-level SW-mapped theory. The Hamiltonian analysis thereby provides three layers of information beyond the Lagrangian diagnosis of Ref. [12]: it localizes the obstruction at the tertiary stage of the Dirac chain, classifies the source dependence through the hierarchical kernel structure $(\Theta^{l},\Sigma^{lk},\Xi^{l})$, and decomposes the on-shell comparison into a Gauss sector recovered exactly and an Ampère sector with irreducible residues.
For generic source profiles with $\Theta^{l}\neq 0$, the tertiary consistency condition feeds back into the primary multiplier and the algorithm closes by multiplier determination rather than by a quaternary constraint. This generic closure is itself a structural result: it establishes how the unrestricted external-current theory terminates canonically within the Banerjee realization. A reduced Hamiltonian description is recovered only after restricting to a sufficient first-class sector in which the displayed source-dependent kernels and the tertiary object become trivial. The gauge generator, reduced brackets, and degree-of-freedom count established in the manuscript are therefore strictly sectorial.
The scope of the paper is correspondingly limited but precise. It does not provide a complete first-/second-class classification for arbitrary external currents, nor a general reduced phase-space formulation of the full theory. Rather, it isolates the source-dependent obstruction in the Dirac chain, clarifies its map-dependent relation to the Euler–Lagrange structure, and identifies the restricted sector in which the standard equal-time Hamiltonian machinery remains available. The extension of this analysis to the minimal current map of Ref. [12] and to the canonical counterpart of the equation-level route are currently under investigation.
Acknowledgements
The authors welcome the support of the Universidad Juárez Autónoma de Tabasco for providing a suitable work environment while this research was carried out. J.M.C. also thanks the Secretariat of Science, Humanities, Technology, and Innovation (Secihti) of México for their support through a grant for postdoctoral studies under Grant No. 3873825. A.G.A.C. acknowledges support from the National Council of Humanities, Science and Technology (CONAHCYT), through a master’s scholarship awarded as part of the national postgraduate funding program.
Statements and Declarations
Competing interests. The authors declare that they have no competing financial or non-financial interests that are directly or indirectly related to the work submitted for publication.
Author contributions. J.M.C. conceived the study, performed the Dirac–Bergmann analysis and the on-shell comparison, and wrote the manuscript. A.G.A.C. developed and implemented the symbolic-algebra toolbox used to verify the displayed formulas, performed computational verification of the results, and contributed to the manuscript preparation. J.M.P.F. supervised the project, contributed to the structural analysis of the constraint chain, and participated in the revision of the manuscript. All authors reviewed and approved the final version.
Data and Code Availability Statement
The results reported here are obtained from analytic and symbolic manipulations; no new experimental data were generated. The Matlab symbolic-algebra toolbox used for the constrained Hamiltonian (Dirac–Bergmann) analysis is available at https://github.com/AlejandroAndarcia/MATLAB-Implementation-of-the-Dirac-Bergmann-Algorithm-for-Field-Theories under the license specified in the repository. Documentation, usage examples, and system requirements are provided there. In the present work the code is used only to verify the analytic formulas displayed in the manuscript; no claim depends exclusively on a hidden computational step.
Appendix A Functional identities for the closure calculation
This appendix collects the functional-derivative identities used in the closure calculation. We only need the contracted combinations that enter the Poisson brackets $\{\phi_{2},H_{\text{ws}}\}$ and $\{\phi_{2},H_{J}\}$. The basic functional derivatives of the secondary constraint are
$\frac{\delta\phi_{2}(x)}{\delta A_{k}(y)}$ $=\frac{1}{2}\,\theta^{kl}\,\partial_{l}^{x}\!\bigl(J^{0}(x)\,\delta^{3}(x-y)\bigr),$ (A.1)
$\frac{\delta\phi_{2}(x)}{\delta\pi^{k}(y)}$ $=\partial_{k}^{x}\,\delta^{3}(x-y).$ (A.2)
Inserting these into the canonical field-theory Poisson bracket gives
$\{\phi_{2},H\}=\int d^{3}y\,\left[\frac{\delta\phi_{2}}{\delta A_{k}(y)}\frac{\delta H}{\delta\pi^{k}(y)}-\frac{\delta\phi_{2}}{\delta\pi^{k}(y)}\frac{\delta H}{\delta A_{k}(y)}\right].$ (A.3)
For the source-free Hamiltonian, only the following contracted combination is needed:
$\{\phi_{2},H_{\text{ws}}\}{}$ $=\frac{1}{2}\,\theta^{kl}(\partial_{l}J^{0})\,\partial_{k}A_{0}$ (A.4)
$-\frac{1}{2}\,\theta^{kl}\eta_{00}\eta_{mk}\partial_{l}(J^{0}\pi^{m}),$
which coincides with Eq. (3.22) in the main text. The result arises from the two terms that generate non-zero contributions at $\mathcal{O}(\theta)$: the linear-in-$\pi$ term $H_{A}=\pi^{k}\partial_{k}A_{0}$ (via the operator $\{\phi_{2}^{(1)},H_{A}\}$) and the quadratic-in-$\pi$ term $H_{D}=-\tfrac{1}{2}\eta_{00}\eta_{ij}\pi^{i}\pi^{j}$ (via $\{\phi_{2}^{(1)},H_{D}\}$). All remaining terms of $H_{\text{ws}}$ contribute zero, either by the Ward gauge identity (terms depending only on $F_{ij}$), by the antisymmetry of $\theta^{kl}$ against symmetric second derivatives, or because they are $\mathcal{O}(\theta^{2})$.
For the source Hamiltonian, the contraction with Eqs. (A.1)–(A.2) yields
$\{\phi_{2},H_{J}\}={}$ $-\partial_{i}J^{i}-\tfrac{1}{2}\theta^{il}(\partial_{i}J^{0})\,\partial_{l}A_{0}$ (A.5)
$+\tfrac{1}{2}\theta^{jl}\partial_{i}\partial_{l}(J^{i}A_{j})-\tfrac{1}{2}\theta^{il}\partial_{i}(J^{k}\partial_{l}A_{k})$
$-\tfrac{1}{2}\theta^{il}\eta_{00}\eta_{lk}\partial_{i}(J^{0}\pi^{k})+\tfrac{1}{2}\theta^{il}\partial_{i}(J^{k}F_{kl}),$
which coincides with Eq. (3.23) of the main text. Each line originates from a specific subset of terms in $\mathcal{H}_{J}$: the first term from $A_{j}J^{j}$; the second from $A_{j}J^{0}\partial_{l}A_{0}$; the third from $A_{j}J^{k}\partial_{l}A_{k}$ (two contributions combined); the fourth from $A_{j}J^{0}\eta_{00}\eta_{li}\pi^{i}$; and the fifth from $A_{j}F_{kl}J^{k}$. The $\pi$-containing terms from $\{\phi_{2},H_{\text{ws}}\}$ (Eq. (A.4)) and from the fourth term here cancel exactly against each other in the sum $\dot{\phi}_{2}$, so that Eq. (3.24) contains no $\pi^{k}$ at all.
Combining Eqs. (A.4) and (A.5) with the explicit derivative (3.20) reproduces Eq. (3.24). The structure of that expression has three key properties. First, all $\pi^{k}$-containing terms cancel algebraically between $\{\phi_{2},H_{\text{ws}}\}$ and $\{\phi_{2},H_{J}\}$, so the constraint surface $\Gamma_{c}=\{\phi_{1}\approx 0,\phi_{2}\approx 0\}$ has no effect on $\dot{\phi}_{2}$. Second, the $\mathcal{O}(\theta^{0})$ piece is $\partial_{\mu}J^{\mu}$: a condition on the external source, not on the phase space, and the direct analogue of the Maxwell consistency condition. Third, the $\mathcal{O}(\theta)$ piece contains $A_{\mu}$ and $\partial_{0}J^{0}$, so the full tertiary candidate depends on the canonical coordinates but not on the canonical momenta. Accordingly, $\{\Phi_{3}^{\rm cand},\Phi_{3}^{\rm cand}\}=0$ whereas the primary–tertiary bracket remains $\{\phi_{1},\Phi_{3}^{\rm cand}\}=-\Theta^{l}\partial_{l}\delta$ in the inhomogeneous-source sector. The next Dirac step therefore feeds back into the primary multiplier through the preservation of $\Phi_{3}^{\rm cand}$, instead of being settled by its self-bracket alone.
Appendix B Constraint brackets through the tertiary stage
This appendix provides the explicit derivations of $\{\phi_{2},\phi_{2}\}$ and $\{\phi_{2},\Phi_{3}^{\rm cand}\}$. These kernels enter the sector discussion of Sec. 3.5.
Secondary self-bracket $\{\phi_{2}(x),\phi_{2}(y)\}$
Using the canonical Poisson bracket in functional form,
$\{F,G\}=\int d^{3}z\,\left[\frac{\delta F}{\delta A_{k}(z)}\frac{\delta G}{\delta\pi^{k}(z)}-\frac{\delta F}{\delta\pi^{k}(z)}\frac{\delta G}{\delta A_{k}(z)}\right],$ (B.1)
and inserting the functional derivatives of $\phi_{2}$ from Eqs. (A.1)–(A.2), one obtains
$\{\phi_{2}(x),\phi_{2}(y)\}={}$ $\int d^{3}z\,\Bigl[\tfrac{1}{2}\theta^{kl}\partial_{l}^{x}\!\bigl(J^{0}(x)\delta^{3}(x-z)\bigr)\,\times$
$\partial_{k}^{z}\delta^{3}(y-z)-\,\partial_{k}^{x}\delta^{3}(x-z)\,\times$
$\tfrac{1}{2}\theta^{kl}\partial_{l}^{z}\!\bigl(J^{0}(y)\delta^{3}(y-z)\bigr)\Bigr].$ (B.2)
Evaluating the first term: since $J^{0}(x)$ is independent of $z$, the factor $\partial_{l}^{x}(J^{0}(x)\delta^{3}(x-z))=(\partial_{l}J^{0}(x))\delta^{3}(x-z)$ when the $x$-derivative acts only on $J^{0}$, plus a total-derivative term that integrates to zero under suitable boundary conditions. After integrating over $z$ using $\delta^{3}(x-z)$,
(first term) $=\tfrac{1}{2}\theta^{kl}(\partial_{l}J^{0}(x))\,\partial_{k}^{y}\delta^{3}(y-x)$ (B.3)
$=-\tfrac{1}{2}\Theta^{k}(x)\,\partial_{k}^{x}\delta^{3}(x-y)\,.$
Evaluating the second term analogously, with the $z$-integration setting $z=y$,
(second term) $=-\tfrac{1}{2}\theta^{kl}(\partial_{l}J^{0}(y))\,\partial_{k}^{x}\delta^{3}(x-y)$ (B.4)
$=\tfrac{1}{2}\Theta^{k}(y)\,\partial_{k}^{y}\delta^{3}(x-y)\,.$
Adding Eqs. (B.3) and (B.4) and using $\partial_{k}^{x}\delta^{3}(x-y)=-\partial_{k}^{y}\delta^{3}(x-y)$,
$\{\phi_{2}(x),\phi_{2}(y)\}{}$ $=\tfrac{1}{2}\Theta^{k}(y)\,\partial_{k}^{y}\delta^{3}(x-y)$ (B.5)
$-\tfrac{1}{2}\Theta^{k}(x)\,\partial_{k}^{x}\delta^{3}(x-y)\,$,
which is manifestly antisymmetric and gives the secondary self-bracket used in Sec. 3.5 of the main text. This bracket vanishes if and only if $\Theta^{k}=0$ everywhere.
Secondary–tertiary bracket $\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}$
Write the compact first-order tertiary candidate as
$\Phi_{3}^{\rm cand}{}$ $=-C+\Theta^{l}\partial_{l}A_{0}+\theta^{li}(\partial_{i}J^{k})\partial_{l}A_{k}$ (B.6)
$-\frac{1}{2}\theta^{il}\partial_{i}\!\bigl(A_{l}C\bigr),\qquad C=\partial_{\mu}J^{\mu},$
where we used the antisymmetry of $\theta^{li}$ to simplify $\theta^{li}\partial_{i}(J^{k}\partial_{l}A_{k})=\theta^{li}(\partial_{i}J^{k})\partial_{l}A_{k}$. Since $\Phi_{3}^{\rm cand}$ contains no canonical momenta and only the $\partial_{m}\pi^{m}$ piece of $\phi_{2}$ has non-vanishing brackets with the spatial gauge potential, one immediately has
$\{\partial_{m}\pi^{m}(x),A_{k}(y)\}$ $=\partial_{k}^{y}\delta^{3}(x-y),$ (B.7)
$\{\partial_{m}\pi^{m}(x),A_{0}(y)\}$ $=0.$
Therefore,
$\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}{}$ $=\,\theta^{li}(\partial_{i}J^{k})(y)\,\partial_{l}^{y}\partial_{k}^{y}\delta^{3}(x-y)$
$-\frac{1}{2}\theta^{il}\partial_{i}^{y}\!\Bigl(C(y)\,\partial_{l}^{y}\delta^{3}(x-y)\Bigr).$ (B.8)
Expanding the last term,
$\theta^{il}\partial_{i}\!\Bigl(C\,\partial_{l}\delta^{3}\Bigr)=\theta^{il}(\partial_{i}C)\partial_{l}\delta^{3}+\theta^{il}C\,\partial_{i}\partial_{l}\delta^{3}.$ (B.9)
The second contribution vanishes because $\partial_{i}\partial_{l}\delta^{3}$ is symmetric in $(i,l)$ whereas $\theta^{il}$ is antisymmetric. One is left with
$\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}{}$ $=\theta^{li}(\partial_{i}J^{k})(y)\,\partial_{l}^{y}\partial_{k}^{y}\delta^{3}(x-y)$ (B.10)
$-\frac{1}{2}\theta^{il}(\partial_{i}C)(y)\,\partial_{l}^{y}\delta^{3}(x-y),$
which is the mixed kernel quoted in Sec. 3.5. The first term depends on spatial current gradients, while the second is driven by gradients of the full divergence $\partial_{\mu}J^{\mu}$.
Kernel decomposition of $\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}$
Equation (B.10) may be decomposed into the symmetric spatial-current kernel and the divergence-gradient kernel introduced in the main text. Defining
$\Sigma^{lk}:=\frac{1}{2}\Bigl(\theta^{li}\partial_{i}J^{k}+\theta^{ki}\partial_{i}J^{l}\Bigr),$ (B.11)
and
$\Xi^{l}:=\theta^{il}\partial_{i}C,$ (B.12)
one obtains
$\{\phi_{2}(x),\Phi_{3}^{\rm cand}(y)\}{}$ $=\Sigma^{lk}(y)\,\partial_{l}^{y}\partial_{k}^{y}\delta^{3}(x-y)$ (B.13)
$-\frac{1}{2}\Xi^{l}(y)\,\partial_{l}^{y}\delta^{3}(x-y),$
because only the symmetric part of the coefficient of $\partial_{l}\partial_{k}\delta^{3}(x-y)$ contributes. This is the form used in Sec. 3.5 to state the explicit kernel conditions for the restricted sufficient first-class subcase.
First-order form of $\dot{\Phi}_{3}^{\rm cand}$
This appendix records the explicit first-order structure of the next Dirac step starting from the tertiary-stage candidate $\Phi_{3}^{\rm cand}=\dot{\phi}_{2}$. Since the candidate contains no momentum variables, its self-bracket vanishes identically. The only bracket relevant for multiplier fixing is the one with the primary constraint,
$\{\Phi_{3}^{\rm cand}(x),\phi_{1}(y)\}$ $=\Theta^{l}(x)\partial_{l}^{x}\delta^{3}(x-y),$ (B.14)
$\Theta^{l}$ $=\theta^{li}\partial_{i}J^{0}.$
Retaining terms through first order in $\theta^{ij}$, one finds the schematic preservation equation quoted in the main text,
$\dot{\Phi}_{3}^{\rm cand}={}$ $\Theta^{l}\partial_{l}u_{1}+\theta^{li}(\partial_{i}J^{m})\partial_{l}\partial_{m}A_{0}$
$-\eta_{00}\theta^{li}(\partial_{i}J_{m})\partial_{l}\pi^{m}+\theta^{li}(\partial_{i}\partial_{0}J^{0})\partial_{l}A_{0}$
$+\theta^{li}(\partial_{i}\partial_{0}J^{k})\partial_{l}A_{k}-\partial_{0}(\partial_{\mu}J^{\mu}),$ (B.15)
where the last term comes from the explicit time dependence of the source-only part of $\Phi_{3}^{\rm cand}$. The first term is the only one involving the arbitrary multiplier. Therefore, in regions where $\Theta^{l}\neq 0$, the condition $\dot{\Phi}_{3}^{\rm cand}\approx 0$ constrains $u_{1}$ along the vector field $\Theta^{l}\partial_{l}$ and closes the generic unrestricted branch by multiplier fixing. Only in degenerate sectors, such as $\Theta^{l}=0$ together with additional cancellations, can the algorithm continue beyond this stage.
The formulas used in the reduced description of Sec. 4 refer only to the further restricted sufficient first-class subcase in which the full tertiary condition $\Phi_{3}^{\rm cand}$ vanishes identically and the equal-time kernels listed in Sec. 3.5 disappear.
Appendix C Independent proof of the divergence identity
This appendix gives an independent derivation of the identity
$\partial_{\nu}\mathcal{E}_{B}^{\nu}=\dot{\phi}_{2},$ (C.1)
using only the primary and secondary constraints, the decomposition (3.19), and the explicit derivative (3.20). No simplified formula from the main text is assumed.
Divergence of the Euler–Lagrange equations
Write the Banerjee-mapped Lagrangian as $\mathcal{L}=\mathcal{L}_{\text{ws}}(F)+\mathcal{L}_{J}$, where
$\mathcal{L}_{J}=-A_{\mu}J^{\mu}-\theta^{\mu\alpha}J^{\beta}A_{\mu}\partial_{\alpha}A_{\beta}+\frac{1}{2}\theta^{\mu\alpha}J^{\beta}A_{\mu}\partial_{\beta}A_{\alpha}.$ (C.2)
Since $\mathcal{L}_{\text{ws}}$ depends only on the antisymmetric field strength, its Euler–Lagrange contribution has identically vanishing divergence. It is therefore enough to vary $\mathcal{L}_{J}$.
A direct application of the Euler–Lagrange operator gives
$\mathcal{E}_{J}^{\nu}={}$ $-J^{\nu}-\theta^{\nu\alpha}J^{\beta}F_{\alpha\beta}-\theta^{\alpha\beta}(\partial_{\alpha}J^{\nu})A_{\beta}$ (C.3)
$\quad-\theta^{\alpha\beta}J^{\nu}\partial_{\alpha}A_{\beta}+\tfrac{1}{2}\theta^{\nu\beta}(\partial_{\mu}J^{\mu})A_{\beta}.$
Taking its divergence and using $\theta^{0i}=0$ yields
$\partial_{\nu}\mathcal{E}_{B}^{\nu}={}$ $-\partial_{\mu}J^{\mu}-\partial_{i}(\theta^{il}J^{\beta}F_{l\beta})-\partial_{\nu}\!\bigl(\theta^{il}(\partial_{i}J^{\nu})A_{l}\bigr)$
$-\partial_{\nu}\!\bigl(\theta^{il}J^{\nu}\partial_{i}A_{l}\bigr)+\frac{1}{2}\partial_{i}\!\bigl(\theta^{il}(\partial_{\mu}J^{\mu})A_{l}\bigr).$ (C.4)
Expanding the time and space derivatives separately, using the antisymmetry of $\theta^{il}$ and the commutativity of partial derivatives, one finds the compact result
$\partial_{\nu}\mathcal{E}_{B}^{\nu}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}+\theta^{li}\partial_{i}(J^{k}\partial_{l}A_{k})$ (C.5)
$\quad-\frac{1}{2}\theta^{il}\partial_{i}\!\bigl[A_{l}(\partial_{\mu}J^{\mu})\bigr].$
Independent computation of $\dot{\phi}_{2}$
Starting from the accepted secondary constraint
$\phi_{2}=\partial_{i}\pi^{i}-J^{0}+\frac{1}{2}\theta^{kl}\partial_{l}(A_{k}J^{0}),$ (C.6)
one has the functional derivatives
$\frac{\delta\phi_{2}(x)}{\delta A_{k}(y)}$ $=\frac{1}{2}\theta^{kl}\partial_{l}^{x}\!\bigl(J^{0}(x)\delta^{3}(x-y)\bigr),$ (C.7)
$\frac{\delta\phi_{2}(x)}{\delta\pi^{k}(y)}$ $=\partial_{k}^{x}\delta^{3}(x-y).$ (C.8)
Inserting them into the functional Poisson-bracket formula and evaluating the contributions of $H_{\text{ws}}$ and $H_{J}$ independently, one obtains
$\{\phi_{2},H_{\text{ws}}\}{}$ $=\frac{1}{2}\theta^{kl}(\partial_{l}J^{0})\partial_{k}A_{0}$
$-\frac{1}{2}\theta^{jl}\eta_{00}\eta_{jn}\partial_{l}(J^{0}\pi^{n}),$ (C.9)
$\{\phi_{2},H_{J}\}={}$ $-\partial_{i}J^{i}-\tfrac{1}{2}\theta^{il}\partial_{i}(J^{0}\partial_{l}A_{0})$
$+\tfrac{1}{2}\theta^{jl}\partial_{i}\partial_{l}(J^{i}A_{j})-\tfrac{1}{2}\theta^{il}\partial_{i}(J^{k}\partial_{l}A_{k})$
$-\tfrac{1}{2}\theta^{il}\eta_{00}\eta_{lk}\partial_{i}(J^{0}\pi^{k})+\tfrac{1}{2}\theta^{il}\partial_{i}(J^{k}F_{kl}).$ (C.10)
The momentum terms cancel exactly in the sum. Adding also the explicit derivative (3.20), one recovers
$\dot{\phi}_{2}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}+\tfrac{1}{2}\theta^{jl}\partial_{i}\partial_{l}(J^{i}A_{j})$
$+\theta^{li}\partial_{i}(J^{k}\partial_{l}A_{k})+\tfrac{1}{2}\theta^{il}\partial_{i}(J^{k}\partial_{k}A_{l})$
$+\tfrac{1}{2}\theta^{kl}\partial_{l}(A_{k}\partial_{0}J^{0}).$ (C.11)
The last three terms may be rearranged algebraically into
$\frac{1}{2}\theta^{il}\partial_{i}\!\bigl[A_{l}(\partial_{\mu}J^{\mu})\bigr].$ (C.12)
so that
$\dot{\phi}_{2}={}$ $-\partial_{\mu}J^{\mu}+\theta^{li}(\partial_{i}J^{0})\partial_{l}A_{0}+\theta^{li}\partial_{i}(J^{k}\partial_{l}A_{k})$
$-\frac{1}{2}\theta^{il}\partial_{i}\!\bigl[A_{l}(\partial_{\mu}J^{\mu})\bigr].$
Comparing Eqs. (C.5) and (C) proves Eq. (C.1). In the unrestricted source problem, the third Dirac-stage candidate is therefore algebraically identical to the canonical pullback of the divergence of the Banerjee Euler–Lagrange equations.